# Mechanical hysterons with tunable interactions of general sign

### v2

@@ Abstract
When an amorphous solid, a disordered magnet, or a crumpled sheet is subjected to an oscillating global drive, it navigates a complex pathway through a multitude of locally-stable states [1]. A powerful approach for understanding these systems has been to focus on the localized hysteretic rearrangements—hysterons—that arise within them [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. In such models, each hysteron $i$ possesses a pair of thresholds $\gamma_{i}^{+}$ and $\gamma_{i}^{-}$ at which it switches between its two possible states, $s_{i}=\pm 1$. More general models incorporate interactions between hysterons [13, 14, 15, 16], so that the switching thresholds $\gamma_{i}^{\pm}$ depart from their “bare” values $\gamma_{i,0}^{\pm}$ as follows:
@@
In addition to offering a foothold for understanding disordered media, Eq. 1 suggests a design principle for a class of synthetic matter. If the thresholds and interactions can be designed, materials could be created with tailored responses to different mechanical environments. As the states $s_{i}$ are readily interpreted as bits, such materials are endowed with digital memory and the ability to perform computations [17]. Several mechanical metamaterials have been developed to this end, including origami bellows [18], corrugated sheets [19], buckled beams [20, 21], and biholar sheets [22, 23]. Yet, even in this realm where one has complete control over the material architecture, no system has been able to realize the full generality of Eq. 1.
Figure 1: Coupled mechanical hysterons with tunable interactions. A. Each hysteron is a rigid bar that may freely rotate between two hard boundaries, under the influence of a driving spring that attaches to a bar applying a horizontal quasistatic global drive, $\gamma$. Hysterons may be coupled together with additional springs. B. Typical force-versus-displacement data for springs used in the experiment. Solid lines are linear fits used to extract the stiffness $k$ and rest length, $x_{0}$. C. Basic behavior of a hysteron. The thresholds, $\gamma^{-}$, $\gamma^{+}$, depend on the geometric properties of the system as well as the state of the other hysterons when interactions are present. D. Experimental realizations with cooperative (left) and frustrated interactions (right).
Here we show that it is possible to embody this abstract hysteron model in a physical system built from simple components: linear springs, rigid rods, and low-friction bearings (Fig. 1). We design interacting mechanical hysterons that allow: (i) tunable switching thresholds for each hysteron, (ii) tunable interactions between the hysterons that can be cooperative or frustrated, as well as reciprocal or non-reciprocal, and (iii) a tunable coupling to a global mechanical drive. We rationalize these results using an exact kinematic model, with a limiting behavior that maps the experimental parameters to the thresholds and interactions of Eq. 1. We then demonstrate the flexibility of the design by realizing a broad set of transition graphs for two interacting hysterons, as well as a “latching” behavior [24] that requires non-reciprocal interactions. Finally, we use larger collections of hysterons to accomplish more complex tasks: distinguishing between even and odd driving cycles [13] and counting the total number of driving cycles.
Our findings provide a general basis for mechanisms that navigate through precise sequences of states, e.g., in robotics [25]. More broadly, our results show that the size of the design space for hysteron-based materials, and the amount of control that can be wielded within it, is perhaps larger than previously thought. Our work thus identifies a vast potential for designed materials that inherently engage with memory [26], as their current state is a by-product of their driving history along with their overall design, which our platform now brings under the experimenter’s control.
@@
Hysteron design.— Figure 1A illustrates the basic bistable unit in our system: a rigid bar that is free to rotate about a central pivot. The bar is corralled between two posts that restrict its angle to an interval, $[\theta^{-},\theta^{+}]$, with $\theta=0$ corresponding to vertical. We achieve precise, low-friction rotation using a radial steel ball bearing with the bearing grease removed. A spring of stiffness $k_{1}$ and rest length $x_{k_{1},0}$ attaches to the bar a distance $L$ from the pivot; the other end of this spring attaches to a long steel shaft at a location $x_{1}$ along its length. The shaft translates freely along its axis via two linear bearings. The $x$-position of this shaft, $\gamma$, will serve as the global mechanical drive. All springs used in the experiments behave linearly, which we verify by measuring their force-versus-displacement curves when pulling up on a known mass resting on an electronic balance (Fig. 1B).
@@
Before investigating multiple coupled rotors, we consider the stability of a single rotor. Starting with the driving rod on the far left and gradually moving it to the right, the rotor begins in the (--) state where $\theta=\theta^{-}$, but becomes unstable and jumps to the ($+$) state ($\theta=\theta^{+}$) when the line running through the spring passes the pivot point of the rod. This transition occurs when the horizontal displacement of the driving rod, $\gamma$, exceeds a threshold $\gamma^{+}=-y\tan\theta^{-}$. Likewise, the rotor flips to ()(-)( - ) when $\gamma<\gamma^{-}=-y\tan\theta^{+}$. The rotor is thus a hysteron: it is monostable outside the interval $[\gamma^{-},\gamma^{+}]$ and it is bistable within (Fig. 1C).
@@
To create interacting hysterons, we space two or more rotors a horizontal distance $w$ apart, and connect each rotor to the driving rod with its own spring of stiffness $k_{i}$ at an anchor point $x_{i}$ on the driving rod, measured from the horizontal position of hysteron $i$ when $\gamma=0$ (Fig. 1D). We may then couple hysterons $i$ and $j$ together with additional springs of stiffness $k_{ij}$ and rest length $x_{0}$, mounted a distance $\ell$ from the pivot point (Fig. 1A). Uncrossed coupling springs create cooperative (ferromagnetic-like) interactions, which encourage two hysterons to occupy the same state, forming the composite states --- - or +++++ + more readily. Crossed springs create frustrated (antiferromagnetic-like) interactions, which encourage them to occupy the states ++-+ - or +-+- + (Fig. 1D). More precisely, the switching thresholds for one hysteron become dependent on the state of the other, giving rise to an expanded set of thresholds. For example, $\gamma^{+}_{1}(s_{2})$ will denote the threshold for hysteron 1 to switch to (+)(+)( + ) when hysteron 2 is in state $s_{2}$. As we will show, these thresholds depend intimately on the geometric parameters in the setup and the properties of the springs.
Figure 2: Transition thresholds for an interacting hysteron pair. A. Using the configuration shown in the left panel of Fig. 1D, we measure the eight transition thresholds as the mounting position of the coupling springs, $\ell$, is varied. The experiments are in good agreement with the theory. Geometric parameters: $y=9.4$ cm, $L=6.3$ cm, $w=15.2$ cm, $x_{1}=0$ cm, $x_{2}=-1.9$ cm, $[\theta_{1}^{-},\theta_{1}^{+}]=[-36.6^{\circ},36.7^{\circ}]$, and $[\theta_{2}^{-},\theta_{2}^{+}]=[-9.75^{\circ},11.2^{\circ}]$. Driving springs: $k_{1}=k_{2}=0.19$ N/cm, rest lengths $3.9$ and $4.2$ cm. Coupling springs: $k_{1,2}=0.07$ N/cm, rest lengths $2.4$ and $2.6$ cm. B. Transition graphs showing behavior for $\ell<4.18$ cm and $4.904.18<\ell<$ 4.90 cm. C. Splitting between pairs of switching thresholds, $\gamma_{i}^{\pm}(+)-\gamma_{i}^{\pm}(-)$, which measures the strength of interactions between hysterons. Triangles: Experiments. Solid lines: Kinematic model. Dashed lines: Eq. 5 for rotors confined to small angles.
@@
Tunable interactions.— The position where the coupling springs attach, $\ell$, offers a convenient way to tune the interaction between the hysterons continuously. To demonstrate this control, we perform experiments on two hysterons with a cooperative interaction, where we fix all other system parameters and vary $\ell$. We measure the switching thresholds by holding one rotor fixed in the (+)(+)( + ) or ()(-)( - ) state, and quasistatically moving the driving rod until the other hysteron undergoes a transition. We measure the rod position with a precision of $0.1$ cm using a ruler affixed under the driving rod. Figure 2A shows the measured values of the eight transitions for this system, at four values of $\ell$.
@@
When $\ell=0$, the spring has no lever arm to exert a torque on either rotor; in practice we remove the coupling springs for these measurements. Here there are only four distinct thresholds, due to the degeneracies $\gamma_{i}^{\pm}(+)=\gamma_{i}^{\pm}(-)$. The particular order of the thresholds, $\gamma^{-}_{1}<\gamma^{-}_{2}<\gamma^{+}_{2}<\gamma^{+}_{1}$, gives rise to a set of stable states and transitions that we may represent in a transition graph [5, 19] – a directed graph that shows the transitions out of each stable state under increasing or decreasing drive (Fig. 2B). For $\ell=1.6$ cm, the thresholds split into eight distinct values. Nevertheless, the same set of stable states and transitions manifest at this interaction strength, as well as for $\ell=3.2$ cm.
@@
Increasing $\ell$ to $4.6$ cm, two pairs of thresholds switch their ordering: $\gamma_{2}^{+}(+)$ falls below $\gamma_{2}^{-}(-)$, and $\gamma_{1}^{+}(+)$ falls below $\gamma_{2}^{+}(-)$. The first switch does not change the transition graph for the system, but the second switch does. Starting in --- -, the first transition upon increasing $\gamma$ is to +-+- + at $\gamma=\gamma_{2}^{+}(-)$. But this state is now unstable, as $\gamma_{2}^{+}(-)>\gamma_{1}^{+}(+)$, meaning that hysteron 1 is above threshold to flip to (+)(+)( + ) as soon as hysteron 2 enters the (+)(+)( + ) state. This event is termed a “vertical avalanche” [15] and is denoted by the arrow from --- - to +++++ + in the second transition graph in Fig. 2B.
@@
Kinematic model.— We can use torque balance to connect the system parameters to the switching thresholds observed in the experiment. First, we assume the system is in a stable state $S$ that specifies the $\{s_{j}\}$. It is then straightforward to write down the Cartesian coordinates of each point where a coupling spring attaches to a rotor in terms of the variables $w$, $\ell$, and $\theta_{j}^{\pm}$. Using the stiffness and rest lengths of the relevant coupling springs, one can then compute the torque $\tau_{ij}(s_{i},s_{j})=\vec{r}_{ij}\crossproduct\vec{F}_{ij}$ of rotor $j$ on rotor $i$. Similarly, one can compute the torque on rotor $i$ by its driving spring, denoted by $\tau_{i}(\gamma,s_{i})$. Hysteron $i$ will undergo a transition when the sum of the torques on it vanish  [^1] To ensure this is an unstable equilibrium, it is necessary that $y>0$, but also that the driving spring $k_{i}$ is sufficiently strong compared to any coupling springs that might disrupt this bistability.:
@@
The $\gamma$ that solve this equation are the switching thresholds, $\gamma_{i}^{\pm}(S)$, wherein hysteron $i$ switches out of the state $s_{i}=\mp 1$  [^2] We do not presently have a precise condition on the parameters that would prohibit additional stable equilibria within the interval $(\theta_{i}^{-},\theta_{i}^{+})$..
@@
Plugging in the experimental parameters for Fig. 2A, we compute these thresholds numerically and compare them with the data. The theory is in good agreement with the experiments. Moreover, the model allows us to anticipate precisely where the qualitative behaviors of the system will change, e.g., the change in the transition graph at $\ell=4.18$ (Fig. 2A,B). Another crossing of switching thresholds occurs at $\ell=4.90$; above this $\ell$ the avalanche from --- - to +++++ + occurs via the intermediate state ++-+ -, as now $\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)$.
Mapping to general model at small rotor angles.— A closer study of Fig. 2A reveals some interesting facts about the system. First, the interactions may be non-reciprocal, meaning that the effect of hysteron 1 on hysteron 2 is not the same as the effect of hysteron 2 on hysteron 1. This can be seen by comparing the splitting of the blue curves for hysteron 2 as compared with the black curves for hysteron 1. Second, the mean of these pairs of thresholds, $[\gamma_{i}^{\pm}(+)+\gamma_{i}^{\pm}(-)]/2$, may shift as a result of the interaction. In principle, one could understand these features of the system by examining the kinematic model, but in practice its nonlinear form conceals an intuitive understanding of these behaviors.
As we now show, we can build such an understanding by studying a limit where the rotor angles are small, $|\theta_{i}^{\pm}|\ll 1$. Moreover, this process will uncover a mapping from the metamaterial parameters to the model parameters of Eq. 1. To emphasize the core relationships in this mapping, we also specialize to the case where the two springs connecting a pair of rotors $i$ and $j$ have identical spring constants $k_{ij}$ and rest lengths $x_{0}\ll w$. We analyze the situation without these restrictions in the Appendix.
We begin by computing the torque of hysteron $j$ on $i$, which under these conditions has a simple form: $\tau_{ij}=2k_{ij}\ell^{2}g_{ij}(\theta_{j}-\theta_{i})$, where the geometric factor $g_{ij}$ is defined by:
@@
We further assume that the rest length $x_{k_{i},0}$ of the driving spring is much less than $L+y$. In this limit, the torque on rotor $i$ by its driving spring is: $\tau_{i}(\gamma,s_{i})=Lk_{i}(\gamma+x_{i}+y\theta_{i})$. We emphasize that this result is for small $\theta_{i}$ but general $\gamma$. To improve the accuracy of the expression at small $\tau_{i}(\gamma,s_{i})$, we replace the term $y\theta_{i}$ with $y\tan\theta_{i}$, which will recover the exact switching thresholds for the case of an uncoupled hysteron.
Figure 3: Cooperative and frustrated interactions. A. We use the same system parameters as in Fig. 2 and calculate switching thresholds from the kinematic model (solid lines, Eq. 2) and the small-$\theta$ model (dashed lines, Eqs. 1, 4 and 5). The small-$\theta$ model captures the gross behavior of the system in spite of the large $\theta_{i}$ involved. B. Same analysis but with crossed springs, producing a frustrated interaction ($J_{ij}<0$) so that now $\gamma_{i}^{\pm}(+)>\gamma_{i}^{\pm}(-)$.
Taking these results together with Eq. 2 and defining $\overline{\theta_{j}}=(\theta_{j}^{+}+\theta_{j}^{-})/2$ and $\Delta\theta_{j}=(\theta_{j}^{+}-\theta_{j}^{-})/2$, we obtain an expression in the form of Eq. 1 with the following correspondences for the bare thresholds and the interactions, respectively:
$\gamma_{i,0}^{\pm}$ $=$ $-x_{i}-y\tan\theta_{i}^{\mp}-2\sum_{j\neq i}\frac{k_{ij}}{k_{i}}\frac{\ell^{2}}{L}g_{ij}(\overline{\theta_{j}}-\theta_{i}^{\mp})$ (4)
$J_{ij}$ $=$ $2\frac{k_{ij}}{k_{i}}\frac{\ell^{2}}{L}g_{ij}\Delta\theta_{j}\ .$ (5)
Note that the sum in Eq. 4 represents the shift of the bare threshold, $\gamma_{i,0}^{\pm}$, due to the presence of interactions, and although it closely resembles the interaction strength $J_{ij}$ in Eq. 5, it is distinct from it. We also note that in this limit, the interactions evidently perturb the $\gamma_{i,0}^{+}$ and the $\gamma_{i,0}^{-}$ by the same amount, i.e., $J_{ij}^{+}=J_{ij}^{-}=J_{ij}$ as in Eq. 1.

### v3

@@ Abstract
When an amorphous solid, a disordered magnet, or a crumpled sheet is subjected to an oscillating global drive, it navigates a complex pathway through a multitude of locally-stable states Paulsen and Keim (2024). A powerful approach for understanding these systems has been to focus on the localized hysteretic rearrangements—hysterons—that arise within them Preisach (1935); Falk and Langer (2011); Keim and Arratia (2014); Mungan et al. (2019); Mungan and Terzi (2019); Terzi and Mungan (2020); Keim et al. (2020); Keim and Medina (2022); Shohat et al. (2023); Muhaxheri and Santangelo (2024); Martínez-Calvo et al. (2024). In these abstract hysteron models, each hysteron $i$ possesses a pair of thresholds $\gamma_{i}^{+}$ and $\gamma_{i}^{-}$ at which it switches between its two possible states, $s_{i}=\pm 1$. More general models incorporate interactions between hysterons Keim and Paulsen (2021); Lindeman and Nagel (2021); van Hecke (2021); Szulc et al. (2022), so that the switching thresholds $\gamma_{i}^{\pm}$ depart from their “bare” values $\gamma_{i,0}^{\pm}$ as follows:
@@
In addition to offering a foothold for understanding disordered media, Eq. 1 suggests a design principle for a new class of synthetic matter. If the thresholds and interactions can be designed, materials could be created with tailored responses to different mechanical environments. As the states $s_{i}$ are readily interpreted as bits, such materials are endowed with digital memory and the ability to perform computations Yasuda et al. (2021). Several mechanical metamaterials have been developed to this end, including origami bellows Jules et al. (2022), corrugated sheets Bense and van Hecke (2021), buckled beams Sirote-Katz et al. (2024); Liu et al. (2024), and biholar sheets Ding and van Hecke (2022); El Elmi and Pasini (2024). Yet, even in this realm where one has complete control over the material architecture, no system has been able to realize the full generality of Eq. 1.
Here, we translate this abstract model into a physical design platform. We show how mechanical hysterons built from rigid bars and linear springs can realize the full generality of Eq. 1 by way of strong, tunable interactions that are ferromagnetic-like or anti-ferromagnetic-like. We then target three specific computations on sequences of mechanical inputs, and build the corresponding systems. Our results provide a replicable, general platform for processing mechanical information in materia.
@@
Hysteron design. Figure 1A illustrates the basic bistable unit in our system: a rigid bar that is free to rotate about a central pivot. The bar is corralled between two posts that restrict its angle to an interval, $\theta^{-},\theta^{+}]$, with $\theta=0$ corresponding to vertical. A spring of stiffness $k_{1}$ and rest length $x_{0}$ attaches to the bar a distance $L$ from the pivot; the other end of this spring attaches to a long steel rod that translates freely along its axis via two linear bearings. The $x$-position of this rod, $\gamma$, serves as the global mechanical drive. All springs used in the experiments behave linearly, which we verify by measuring their force-versus-displacement curves when pulling up on a known mass resting on an electronic balance (Fig. 1B). We analyze these data to measure $k$ and $x_{0}$ for each spring used in the experiment.
@@
We first consider the stability of a single rotor. Starting with the driving rod on the far left and gradually moving it to the right, the rotor begins in the (-) state where $\theta=\theta^{-}$, but becomes unstable and jumps to the (++) state ($\theta=\theta^{+}$) when the line running through the spring passes the pivot point of the rod. This transition occurs when the horizontal displacement of the driving rod, $\gamma$, exceeds a threshold $\gamma^{+}=-y\tan\theta^{-}$. Likewise, the rotor flips to ()(-) when $\gamma<\gamma^{-}=-y\tan\theta^{+}$. The rotor is thus a hysteron: it is monostable outside the interval $\gamma^{-},\gamma^{+}]$ and it is bistable within (Fig. 1C). We may move this interval of stability by shifting the position $x_{i}$ where hysteron $i$ couples to the driving rod (where $x_{i}=0$ corresponds to no shift). Supplementary Movie 1 shows these basic behaviors for a pair of uncoupled rotors.
@@
Next, we create interactions by coupling rotors $i$ and $j$ together with additional springs of stiffness $k_{ij}$ and rest length $x_{0}$, mounted at a distance $\ell$ from the pivot point (Fig. 1A). Uncrossed coupling springs create cooperative (ferromagnetic-like) interactions, which encourage two hysterons to occupy the same state, forming the composite states -- or ++++ more readily. Crossed springs create frustrated (antiferromagnetic-like) interactions, which encourage them to occupy the states ++- or +-+ (Fig. 1D).
These interactions may be understood by torque balance. Suppose two rotors are held fixed at $\theta_{1}=\theta_{2}=0^{\circ}$. Coupling the rotors with two identical springs leads to zero net torque on rotor $1$, as the torques from the two springs cancel whether they are crossed or uncrossed. Moving rotor $2$ into the ++ state increases the distance spanned by one coupling spring and decreases the distance spanned by the other. Since these springs attach to rotor $1$ on different sides of the pivot, this change in the torque adds up—two positive numbers for uncrossed springs (pushing rotor 1 towards the ++ state) and two negative values for crossed springs (pushing rotor 1 towards the - state). Through these interactions, the switching thresholds for one hysteron thus become dependent on the state of the other, giving rise to an expanded set of thresholds. The ensuing complexity and rapid tunability is demonstrated in Supplementary Movie 2, where we demonstrate a multitude of transitions for two cooperatively-coupled hysterons.
@@
Tunable interactions. The position where the coupling springs attach, $\ell$, offers a convenient way to tune the interaction between the hysterons continuously. To demonstrate this control, we perform experiments on two hysterons with a cooperative interaction, where we fix all other system parameters and vary $\ell$. We measure the switching thresholds by holding one rotor fixed in the $(+)$ or ()(-) state, and quasistatically moving the driving rod until the other hysteron undergoes a transition. For example, $\gamma^{+}_{1}(s_{2})$ denotes the measured threshold for hysteron 1 to switch to $(+)$ when hysteron 2 is in state $s_{2}$. Figure 2A shows the measured values of the eight transitions for this system, at four values of $\ell$.
@@
When $\ell=0$, the coupling springs have no lever arm to exert a torque on either rotor  [^1] In practice we remove the coupling springs for the case $\ell=0$.. Here there are only four distinct thresholds, due to the degeneracies $\gamma_{i}^{\pm}(+)=\gamma_{i}^{\pm}(-)$. The particular order of the thresholds, $\gamma^{-}_{1}<\gamma^{-}_{2}<\gamma^{+}_{2}<\gamma^{+}_{1}$, gives rise to a set of stable states and transitions that we may represent in a transition graph Mungan et al. (2019); Bense and van Hecke (2021) – a directed graph that shows the transitions out of each stable state under increasing or decreasing drive (Fig. 2B). For $\ell=1.6$ cm, the thresholds split into eight distinct values. Nevertheless, the same stable states and transitions remain at $\ell=1.6$ cm and at $\ell=3.2$ cm.
@@
Increasing $\ell$ to $4.6$ cm, two pairs of thresholds switch their ordering: $\gamma_{2}^{+}(+)$ falls below $\gamma_{2}^{-}(-)$, and $\gamma_{1}^{+}(+)$ falls below $\gamma_{2}^{+}(-)$. The first switch does not change the transition graph for the system, but the second does. Starting in --, the first transition upon increasing $\gamma$ is to +-+ at $\gamma=\gamma_{2}^{+}(-)$. But this state is now unstable, as $\gamma_{2}^{+}(-)>\gamma_{1}^{+}(+)$, meaning that hysteron 1 is above threshold to flip to $(+)$ as soon as hysteron 2 enters the $(+)$ state. This “vertical avalanche” van Hecke (2021) is denoted by the arrow from -- to ++++ in the second transition graph in Fig. 2B.
This tunability of the thresholds and interactions suggests that our mechanical hysterons should be able to exhibit a wide variety of behaviors. We confirm this notion in Supplementary Movies 2-5, where we demonstrate all possible transition graphs for two hysterons with cooperative interactions ($J_{ij}>0$) and frustrated interactions ($J_{ij}<0$) van Hecke (2021). This suite of behaviors includes a 2-bit analog-to-digital converter (Fig. S1A, transition graph $ix$). We provide further descriptions of these experiments in the Supplementary Information and in Fig. S1.
@@
Kinematic model. We can use torque balance to connect the system parameters to the switching thresholds observed in the experiment. First, we assume the system is in a stable state $S$ that specifies the $\{s_{j}\}$. It is then straightforward to write down the torque $\tau_{ij}(s_{i},s_{j})$ of rotor $j$ on rotor $i$, in terms of the properties and mounting positions of the coupling springs. Similarly, one can compute the torque $\tau_{i}(\gamma,s_{i})$ on rotor $i$ by its driving spring. Exact expressions for $\tau_{i}(\gamma,s_{i})$ and $\tau_{ij}(s_{i},s_{j})$, which take into account the finite rest length of the springs, are provided in the Supplementary Information. Hysteron $i$ will undergo a transition when the sum of the torques on it vanish  [^2] To ensure this is an unstable equilibrium, it is necessary that $y>0$, but also that the driving spring $k_{i}$ is sufficiently strong compared to any coupling springs that might disrupt this bistability.:
@@
The $\gamma$ that solve this equation are the switching thresholds, $\gamma_{i}^{\pm}(S)$, wherein hysteron $i$ switches out of the state $s_{i}=\mp 1$  [^3] We do not presently have a precise condition on the parameters that would prohibit additional stable equilibria within the interval $(\theta_{i}^{-},\theta_{i}^{+})$.. We use Eq. 2 to compute the predicted thresholds for Fig. 2A and find good agreement with the experiments. Moreover, the model allows us to anticipate precisely where the qualitative behaviors of the system will change, e.g., the change in the transition graph at $\ell=4.18$ (Fig. 2A,B)  [^4] Another crossing of switching thresholds occurs at $\ell=4.90$; above this $\ell$ the avalanche from -- to ++++ occurs via the intermediate state ++-, as now $\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)$..
Linear interactions at arbitrary rotor angles. Equation 1 posits that the hysteron interactions are pairwise and linear. We now show that our system exhibits such interactions when the rest length $x_{0}$ of each driving spring is much less than $L+y$. In this limit, the torque on rotor $i$ by its driving spring is: $\tau_{i}=Lk_{i}\cos\theta_{i}(\gamma+x_{i}+y\tan\theta_{i})$. We emphasize that this result is for arbitrary $\theta_{i}$ and $\gamma$ as detailed in the Supplementary Information. Plugging this expression into Eq. 2 and solving for $\gamma$ gives the transitions $\gamma^{\pm}_{i}(S)$ for rotor $i$ when the system is in microstate $S$. The interaction between hysterons $i$ and $j$ may then be calculated by the difference, $\gamma_{i}^{\pm}(S|_{s_{j}=-})-\gamma_{i}^{\pm}(S|_{s_{j}=+})=2J_{ij}$, which yields:
$J_{ij}^{\pm}=\frac{1}{2Lk_{i}\cos\theta_{i}^{\mp}}(\tau_{ij}(\mp,+)-\tau_{ij}(\mp,-))\ ,$ (3)
independent of the state of all other rotors. Such a result establishes that the interactions are pairwise and linear, even for arbitrary rotor angles and finite coupling spring rest lengths. This linearity may be attributed to two key features of our system: (i) the affine relationship between the drive $\gamma$ and its torque on a rotor, and (ii) the additivity of the torques from the coupling springs.
Equation 3 additionally says that there are two sets of interactions strengths: $J_{ij}^{+}$ describes the interactions for hysteron $i$ when $\theta_{i}=\theta_{i}^{-}$, whereas $J_{ij}^{-}$ is the relevant matrix when $\theta_{i}=\theta_{i}^{+}$. Such abstract hysteron models have been proposed recently Szulc et al. (2022), where $J_{ij}$ in Eq. 1 is replaced by $J_{ij}^{\pm}$. In Fig. S2 we show that the simpler scenario, $J_{ij}^{+}=J_{ij}^{-}=J_{ij}$, occurs when the stopping posts are positioned symmetrically for each rotor, meaning $\theta_{i}^{+}=-\theta_{i}^{-}$ for each $i$.
We can also calculate the bare thresholds, $\gamma_{i,0}^{\pm}$, in terms of the torques on rotor $i$. They are found from the relation, $\gamma^{\pm}_{i}(S^{+})+\gamma^{\pm}_{i}(S^{-})=2\gamma_{i,0}^{\pm}$, where $S^{\pm}$ is the state with $s_{j}=\pm 1$ for all $j\neq i$. Using Eq. 2 and the expression for $\tau_{i}$ above yields:
$\gamma_{i,0}^{\pm}=-x_{i}-y\tan\theta_{i}^{\mp}-\frac{1}{2Lk_{i}\cos\theta_{i}^{\mp}}\sum_{j\neq i}(\tau_{ij}(\mp,+)+\tau_{ij}(\mp,-))\ .$ (4)
We emphasize that the sum in Eq. 4 represents a shift of the bare threshold due to the presence of interactions. Physically, it arises because the torque exerted by rotor $j$ on rotor $i$, averaged over the two states $s_{j}=\pm 1$, may be non-zero. By contrast, the interactions $J_{ij}$ measure the positive and negative departures from this mean, which are then added on top of this bare threshold, following Eq. 1.
Mapping to hysteron model at small rotor angles. We can gain a deeper understanding of the interactions by studying a limit where the rotor angles are small, $|\theta_{i}^{\pm}|\ll 1$. To emphasize the core relationships in this mapping, we specialize to the case where the two springs connecting rotors $i$ and $j$ have identical spring constants $k_{ij}$ and rest lengths $x_{0}$. To simplify expressions, we consider coupling springs with rest lengths that are small compared to either the rotor spacing or the mounting distance on the rotor: $x_{0}\ll w$ or $x_{0}\ll\ell$. We show in the Supplementary Information that under these conditions, the torque of hysteron $j$ on $i$ has a simple form: $\tau_{ij}=2g_{ij}k_{ij}\ell^{2}(\theta_{j}-\theta_{i})$, where the geometric factor $g_{ij}$ is defined by:
@@
Plugging this torque into Eqs. 3 and 4 at small angles, we obtain:
$J_{ij}$ $=$ $g_{ij}\frac{k_{ij}}{k_{i}}\frac{\ell^{2}}{L}(\theta_{j}^{+}-\theta_{j}^{-})\ ,$ (6)
$\gamma_{i,0}^{\pm}$ $=$ $-x_{i}-y\tan\theta_{i}^{\mp}-\sum_{j\neq i}\frac{k_{ij}}{k_{i}}\frac{\ell^{2}}{L}g_{ij}(\theta_{j}^{+}+\theta_{j}^{-}-2\theta_{i}^{\mp})\ ,$ (7)
where we retained $y\tan\theta_{i}^{\mp}$ in place of $y\theta_{i}^{\mp}$ in Eq. 7 so that it gives the exact threshold in the absence of coupling.
Figure S3 shows that the small-$\theta$ model of Eqs. 1, 6, and 7 provides a good approximation to the exact kinematic model, Eq. 2, for both cooperative and frustrated interactions. Moreover, Eqs. 6 and 7 provide important physical insights into our system. For instance, they tell us that only one interaction matrix is sufficient to describe the system at small rotor angles ($J_{ij}^{+}=J_{ij}^{-}=J_{ij}$), even without the additional symmetry of the rotor angles assumed in Fig. S2. They also identify an inherent non-reciprocity of the interactions that we explore in the next section. More valuable still, Eqs. 6 and 7 are a blueprint for designing the thresholds and interactions in our system at will—an enticing topic that comprises the remainder of this article.